### v2

# Quantifying the dynamic structural resilience of international staple food trade networks: An entropy-based approach

Abstract
Establishing a resilient food trade system is an international consensus on safeguarding food security amid growing disruptions. We employ an entropy-based approach to quantify the dynamic structural resilience of international trade networks for maize, rice, soybean, and wheat from 1986 to 2022. Using index decomposition analysis, we also investigate the relative contributions of several internal components to changes in resilience. Within this framework, despite heterogeneity across different food commodities, we find that current trade networks are relatively redundant, with improvements in efficiency being the dominant driver of changes in resilience. In addition, we reveal a historically pronounced impact of flow concentrations on resilience, while trade interactions have become increasingly important in recent years. Furthermore, following the leave-one-out approach, we identify critical economies and trade relationships that disproportionately affect the overall resilience, many of which are overlooked by conventional volume-based or centrality-based metrics. Moreover, we highlight that the overconcentration of flows along core trade relationships may undermine the overall efficiency and resilience, whereas peripheral trade networks play an essential role in sustaining resilience, underscoring the importance of promoting more equitable trade relations. These findings not only provide new insights for assessing the resilience of international food trade systems but also propose directions for making them more resilient.
keywords:
International food trade network; Food security; Network resilience; Leave-one-out approach; Index decomposition analysis
articletype: RESEARCH ARTICLE
1 Introduction
Food security, an increasingly prominent global concern, was traditionally conceptualized as comprising three core pillars: availability, access, and utilization (Barrett, 2010). Recently, this framework has been extended to incorporate three additional dimensions—stability, agency, and sustainability—thus forming a comprehensive six-dimensional paradigm (Clapp et al., 2022). Historically, food insecurity has been driven primarily by shortages in food availability, but in modern times, global famine and destitution are primarily caused by limited access to food (Suweis et al., 2015). Pandemics, extreme climate changes, pests, and geopolitical conflicts significantly destabilize food availability and access (Naqvi et al., 2020; Nicholson et al., 2021). Stabilizing these two pillars is therefore a key component of contemporary food security strategies (Barrett, 2010).
Within the framework of food security, international food trade contributes to both food availability and accessibility (Traverso and Schiavo, 2020), although its role in safeguarding food security remains a subject of debate (Farsund et al., 2015; Grassia et al., 2022). Some scholars emphasize the positive contribution of food trade to food security (Dorosh et al., 2009; Grassia et al., 2022). For example, Dithmer and Abdulai (2017) argue that food trade can help mitigate domestic excess demand or supply, thereby providing countries with more options to enhance food availability than under a self-sufficiency policy, and contributing to greater stability in national food supply and prices. Similarly, Traverso and Schiavo (2020) find that participation in international food trade positively affects aggregate food availability and access in low-income countries, from a perspective of macronutrient flows. In contrast, other scholars highlight its potential downsides (Headey, 2011; Odey et al., 2023). For instance, Bouët et al. (2005) argue that the benefits of trade may be overestimated, particularly for developing countries lacking a comparative advantage in agricultural production. Based on evidence from developing economies, Mary (2019) suggests that pursuing food self-sufficiency is more beneficial in the short term, even if such policies conflict with World Trade Organization regulations and the prevailing global trade agenda.
Although the role of international trade in food security remains contested, there is broad consensus on the necessity of establishing a resilient international food trade system (Tu et al., 2019; FAO, 2021; Engemann et al., 2023). Achieving this goal, however, requires addressing two foundational questions: (1) what constitutes resilience in the context of the international food trade system? and (2) how can it be effectively measured?
Resilience is a fundamental property of systems (Liu et al., 2022), yet it remains a contested construct, with ongoing and often discipline-specific debates among various schools of thought, including those in engineering, psychology, management, and economics (Youssef and Luthans, 2007; Francis and Bekera, 2014; Brunnermeier, 2024; Padovano and Ivanov, 2025; Gunasekaran et al., 2015; Ivanov, 2024). Its concept originates in ecology, where Holling (1973) defines it as a system’s capacity to withstand disturbances without collapsing or transitioning to a fundamentally altered state, and to recover its essential functions afterward. Building on this concept, scholars initially define food system resilience as the ability to maintain a stable food supply—both in terms of quantity and diversity—while withstanding disruptions (Tendall et al., 2015). As the food trade system, accounting for approximately 20% of the food consumed globally, constitutes a critical subsystem of the broader food system (D’Odorico et al., 2014), its resilience is often conceptualized in a similar manner (Mena et al., 2022; Jafari et al., 2024).
While efforts to accurately define the resilience of the food trade system continue, scholars also initiate discussions on how it can be effectively measured, yet no consensus emerges (Pimm et al., 2019). Most resilience measures are based on the concept of the resilience triangle (see Fig. 1 in Bruneau et al. (2003)), which defines resilience as the level of system performance retained during a disturbance period (Cheng et al., 2022). For example, studies in engineering often assess resilience by measuring the cumulative loss in infrastructure quality over time (Fang et al., 2022; Chen et al., 2024; Sahebjamnia et al., 2015). In economics, resilience is often understood as an economic system’s dynamic response to external shocks, and can be thus measured through response functions (Klimek et al., 2019; Brunnermeier, 2024). In ecology, resilience is typically quantified in terms of recovery rate or the time required to adapt to changing disturbance regimes (Seidl et al., 2016). Accordingly, Wei et al. (2025) extend this concept to “V-shaped curve” and assess resilience of food trade networks from a port perspective. These studies emphasize the system’s speed of recovery and the magnitude of loss, offering valuable insights into resilience measurement. However, these approaches may be inadequate, as they focus primarily on system performance following a disturbance while neglecting that in the absence of shocks, which may not completely reflect the fact that resilience is a fundamental system property (Liu et al., 2022). Meanwhile, determining the duration of shocks and the appropriate recovery level remains challenging and unrealistic (Hosseini et al., 2016).
Most real-world systems are structured as networks with complex topologies (Barabási and Albert, 1999). Drawing on approaches from network science, a substantial body of literature assesses system resilience by analyzing topological metrics of the network. For example, resilience can be assessed using the inverse of the harmonic mean of the shortest path lengths between nodes, referred to as global efficiency (which differs from the notion of efficiency used in our study) (Latora and Marchiori, 2001; Osei-Asamoah and Lownes, 2014; Testa et al., 2015; Xie et al., 2021; Wei et al., 2022; Chen et al., 2024), which is also applied to assess resilience in the global agri-food wheat trade system (Fair et al., 2017). Moreover, Dixit et al. (2020) measure supply chain resilience by integrating four topological properties, including density, centrality, size, and connectivity, which is also applied to assess resilience in crude oil and gas trade networks (Mou et al., 2020b; Sun et al., 2023). Although topological properties primarily capture the structural aspects of a networked system, they overlook its dynamic characteristics (Qi and Mei, 2024) and do not adequately reflect the actual dynamic evolution of system resilience, particularly in food trade system that is inherently directed and weighted. Notably, within a framework proposed by Gao et al. (2016), existing literature can infer whether a system is resilient based on dynamic equations (Xu et al., 2019; Zhao et al., 2024; Tu et al., 2019), which is also applied to food trade systems (Wei et al., 2025). However, this resilience inference has to rely on the strong assumptions that the degrees of interconnected nodes are largely independent (Liu et al., 2024a). These assumptions contrast with the high assortativity observed in international trade networks (Mou et al., 2020a), which may limit the applicability of such framework in this context.
Information theory is introduced to fill the gaps in measuring resilience of trade systems. Although resilience encompasses several principles such as adaptability and cohesion (Bhamra et al., 2011), diversification is widely regarded as the most fundamental, as it manifests at multiple levels and across different components of the food system (Hertel et al., 2021). Higher diversification implies a broader distribution of probabilities, indicating that the food system is less likely to collapse in the face of disturbances (Nicholson et al., 2021). In this context, entropy, an information-theoretic measure of uncertainty and diversity, provides a more effective representation of system resilience, particularly for systems characterized by directed and weighted flows. Among related studies, entropy is first applied to assess the stability of food webs in ecosystems (Rutledge et al., 1976; Ulanowicz, 1979). Specifically, based on the entropy-based approach, the trade-off between two underlying system properties—ascendency and reserve—is proposed to measure resilience of ecosystems (Ulanowicz, 2009; Ulanowicz et al., 2009). Kharrazi et al. (2017) argue that flows of energy in food webs are inherently exchanges of resources and subsequently assess resilience of global commodity trade systems by redefining the above two properties as efficiency and redundancy. Building on this line of research, scholars recently begin to apply this approach to measure resilience in various complex systems, such as phosphorus cycling networks (Liang et al., 2020), coupled nitrigen-phosphorus cycling networks (Luo et al., 2024), waste metals trade systems (Zisopoulos et al., 2025), and power systems (Panyam et al., 2019), offering interesting insights for our understanding.
Overall, there are substantial academic discussions on the definitions and metrics of resilience, with many of these concepts applied to food systems. However, in our view, research specifically addressing the resilience of food trade systems remains underdeveloped, particularly in the context of increasing geopolitical conflicts and climate risks nowadays. First, most existing studies equate resilience with some unweighted topological properties of the food trade system, which may overlook its dynamic characteristics and the influence of trade volume on overall resilience. Few studies employ information-theoretic entropy to measure the resilience of food trade systems (Li et al., 2024), highlighting a critical gap that warrants further attention. Second, existing studies primarily examine the correlations between the resilience of food trade systems and determinants such as topological properties and external shocks (Li et al., 2024; Ji et al., 2024). However, few provide recommendations for enhancing resilience based on an analysis of its internal components derived through decomposition methods. Such decomposition facilitates a detailed attribution of system resilience to its structural and functional components, thereby revealing its driving mechanisms. Third, although some studies investigate the importance of individual economies within the trading system (Grassia et al., 2022), most rely on indicators such as trade volume and centrality (Gutiérrez-Moya et al., 2021; Zhang and Zhou, 2023), potentially overlooking important nodes characterized by moderate trade volumes and fewer trading partners but playing pivotal structural roles. Moreover, link-level disruptions are more frequent in real-world trade systems, yet existing research predominantly focuses on the significance of nodes, while the importance of trade links remains relatively underexplored (Qian et al., 2017; Guo et al., 2025). Thus, assessing their importance is necessary, especially from a resilience perspective.
To address these gaps and concerns, the following research questions frame our work:
  1. RQ1:
    How can the resilience of the international food trade system be effectively measured?
  2. RQ2:
    How do internal components contribute to changes in resilience of the international food trade system?
  3. RQ3:
    How can critical economies and trade relationships in the international food trade system be identified from a resilience perspective?
The subsequent sections of this paper are structured in the following manner. The methodological framework is presented in Section 2. Section 3 describes the international food trade network and its resilience. Then, Section 4 and Section 5 present and discuss the main findings. Finally, Section 6 serves as the conclusion, which includes the contributions, implications, limitations, and future outlook.
2 Methodology
2.1 Topological properties of the international food trade system
The international food trade system is usually regarded as a network that exhibits small-world (e.g., high clustering coefficients coupled with short average path lengths) and scale-free (e.g., power-law degree distributions) topological properties in economies’ connectivity (Fagiolo et al., 2009). A food trade network is a weighted digraph $G=\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V}$ represents the vertex set with $N_{\mathcal{V}}$ economies and $\mathcal{E}$ represents the link set with $N_{\mathcal{E}}$ trade relationships. $f_{ij}$ denotes the volume of trade flow that economy $i$ exports to economy $j$. There are several commonly used topological metrics for characterizing network structure, which are briefly described below.
  1. (1)
    The in-degree $k_{i}^{\mathrm{in}}$ and out-degree $k_{i}^{\mathrm{out}}$ of economy $i$ are defined respectively as the numbers of economies from which it imports and to which it exports.
  2. (2)
    The in-strength $s_{i}^{\mathrm{in}}=\sum\limits_{i=1}^{N_{\mathcal{V}}}f_{ij}$ and out-strength $s_{i}^{\mathrm{out}}=\sum\limits_{j=1}^{N_{\mathcal{V}}}f_{ij}$ of economy $i$ are defined as the total volumes of imports and exports, respectively.
  3. (3)
    The average degree and average strength are respectively calculated as: $\langle k\rangle=\frac{1}{N_{\mathcal{V}}}\sum\limits_{i=1}^{N_{\mathcal{V}}}k_{i}$ and $\langle s\rangle=\frac{1}{N_{\mathcal{V}}}\sum\limits_{i=1}^{N_{\mathcal{V}}}s_{i}$, where $k_{i}=k_{i}^{\mathrm{in}}+k_{i}^{\mathrm{out}}$ and $s_{i}=s_{i}^{\mathrm{in}}+s_{i}^{\mathrm{out}}$.
  4. (4)
    The average clustering coefficient represents the average level of interconnectedness among the trade partners of economies, defined as $\langle C\rangle=\frac{1}{N_{\mathcal{V}}}\sum\limits_{i=1}^{N_{\mathcal{V}}}C_{i}$, where the local clustering coefficient of $i$ is $C_{i}=\frac{T_{i}}{k_{i}^{\mathrm{in}}(k_{i}^{\mathrm{in}}-1)+k_{i}^{\mathrm{out}}(k_{i}^{\mathrm{out}}-1)}$ and $T_{i}$ represents the number of actual connections among the trade partners of economy $i$.
  5. (5)
    The density represents the level of trade closeness among economies, defined as: $\rho=\frac{N_{\mathcal{E}}}{N_{\mathcal{E}}^{\max}}=\frac{N_{\mathcal{E}}}{N_{\mathcal{V}}(N_{\mathcal{V}}-1)}$, where $N_{\mathcal{E}}^{\max}$ represents the maximum number of links possible in the network.
  6. (6)
    The average shortest path length represents the average level of separation among economies, defined as $L=\frac{1}{N_{\mathcal{V}}(N_{\mathcal{V}}-1)}\sum\limits_{i\neq j\in\mathcal{V}}d_{ij}$, where $d_{ij}$ represents the length of the shortest path connecting economy $i$ to economy $j$.
  7. (7)
    The diameter represents the length of the longest $d_{ij}$, defined as $D=\max\limits_{i\neq j\in\mathcal{V}}d_{ij}$.
  8. (8)
    Modularity represents the effectiveness of partitioning a network into communities, defined as $Q=\frac{1}{2N_{\mathcal{V}}\langle s\rangle}\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\left(f_{ij}-\frac{s_{i}s_{j}}{2N_{\mathcal{V}}\langle s\rangle}\right)\delta(c_{i},c_{j})$, where $\delta(c_{i},c_{j})=1$ if economies $i$ and $j$ belong to the same community and $\delta(c_{i},c_{j})=0$ otherwise. The Louvain algorithm, recognized as one of the fastest and best performing algorithms in comparative analysis (Blondel et al., 2008; Lancichinetti and Fortunato, 2009), is employed for community detection in this paper.
2.2 Efficiency and redundancy of the international food trade system
In information theory, entropy measures a system’s uncertainty (or diversity). A higher entropy indicates a greater unpredictability and a higher diversity in this system’s states. Accordingly, we can adopt joint entropy to characterize a system’s overall structural properties (Rutledge et al., 1976; Ulanowicz, 1979), which is also called as “capacity” for system development (Ulanowicz and Norden, 1990). Given that joint entropy is equal to the sum of mutual information and conditional entropy, we further introduce two underlying quantities, efficiency and redundancy, which demonstrate opposing properties of a system. Efficiency embodies the flow articulation or constraints within networked configurations, which tends to increase due to preferential interactions between economies (Kharrazi et al., 2017). As economies become more interdependent, trade flows tend to become more concentrated and specialized, thereby enhancing overall system efficiency. In information theory, a higher mutual information indicates a stronger statistical inter-dependence between variables. Let $s=\sum\limits_{i=1}^{N_{\mathcal{V}}}s_{i}^{\mathrm{out}}=\sum\limits_{j=1}^{N_{\mathcal{V}}}s_{j}^{\mathrm{in}}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}f_{ij}$, the efficiency $e$ can be expressed as:
$e=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{f_{ij}}{s}\ln\frac{f_{ij}s}{s_{i}^{\mathrm{out}}s_{j}^{\mathrm{in}}},$ (1)
which is also called as the average mutual information (Rutledge et al., 1976; Ulanowicz and Norden, 1990), where
$p_{ij}=\frac{f_{ij}}{s}$ (2)
denotes the concentration degree of trade flows from economy $i$ to $j$. $p_{i}^{\mathrm{out}}$ and $p_{j}^{\mathrm{in}}$ respectively denote the concentration degree of trade outflows from economy $i$ to the others:
$p_{i}^{\mathrm{out}}=\frac{s_{i}^{\mathrm{out}}}{s},$ (3)
and that of trade inflows from the others to economy $j$:
$p_{j}^{\mathrm{in}}=\frac{s_{j}^{\mathrm{in}}}{s}.$ (4)
Redundancy embodies the diversity of food trade pathways, which is critical for the system’s capacity adapting to changing environmental conditions arising from shocks or disturbances (Luo et al., 2024). A greater number of pathways in food trade allows economies to have greater freedom in the trade market and enhances the system’s robustness in coping with risks. In information theory, conditional entropy measures the uncertainty remaining given that some information is known. A higher conditional entropy indicates a greater availability of alternative pathways in food trade and a consequently higher system redundancy. Hence, the redundancy $r$ can be expressed as
$r&=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln p_{j\mid i}-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln p_{i\mid j}\\ &=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}}-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{j}^{\mathrm{in}}}\\ &=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{f_{ij}}{s}\ln\frac{s_{i}^{\mathrm{out}}s_{j}^{\mathrm{in}}}{f_{ij}^{2}}\end{aligned}.$ (5)
2.3 Resilience of the international food trade system
Based on these two system properties, the ratio $\alpha$, a more comprehensive metric to indicate the order of a system, is proposed for reflecting the trade-off between efficiency and redundancy in the food trade system, which is expressed as
$\alpha=\frac{e}{e+r},$ (6)
where $0\leq\alpha\leq 1$. Inspired by ecological systems that their orders are all close to $1/{\mathrm{e}}$ (Zorach and Ulanowicz, 2003), Ulanowicz (2009) defined the fitness $R$ of a system for change to be the product of $\alpha$ and the Boltzmann measure of its disorder such that $R=-c\alpha\ln\alpha$, where $c$ is an appropriate scalar constant. When $\alpha=1/{\mathrm{e}}$, $R^{\prime}=0$, which means the order of the system is optimized. The underlying assumption is that ecosystems exhibit superior trade-off because they have undergone long-term natural selection (Ulanowicz, 2009; Liang et al., 2020; Kharrazi et al., 2013). Subsequently, let $c=1$, the resilience $R$ of a food trade network can be defined as
$R=-\alpha\ln\alpha,$ (7)
where the maximum value is $1/{\mathrm{e}}\approx 0.3679$. Hence, a resilient food trade system possesses both the capacity to flexibly allocate food (efficient) and to absorb shocks via several pathways (redundant). When $\alpha>1/{\mathrm{e}}$, the system is more efficient and productive but more vulnerable, and vice versa it is more redundant but more inefficient. Notably, resilience vanishes when $\alpha=0$ (overly redundant) or $\alpha=1$ (overly efficient).
2.4 Relative contributions of changes in efficiency and redundancy
Given that resilience is constituted by efficiency and redundancy, we can employ the index decomposition analysis to investigate the relative contributions of changes in these two properties to changes in resilience (Ang and Zhang, 2000; Liang et al., 2020). We can rewrite Eq. (7) as
$R=-\frac{e}{e+r}\ln\left(\frac{e}{e+r}\right).$ (8)
Changes in resilience with time $t$ can be defined as
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{1+\ln(e)-\ln(e+r)}{\left(e+r\right)^{2}}\left(e\frac{\mathrm{d}r}{\mathrm{d}t}-r\frac{\mathrm{d}e}{\mathrm{d}t}\right).$ (9)
As we can use $\Delta R$, $\Delta e$, and $\Delta r$ to substitute $\mathrm{d}R$, $\mathrm{d}e$, and $\mathrm{d}r$, respectively, Eq. (9) can be rewritten as
[…]
where $\Delta_{e}R$ and $\Delta_{r}R$ refer to the contributions of changes in efficiency and redundancy to changes in resilience, respectively.
Subsequently, we define a ratio as follows (Liang et al., 2020):
$\mathrm{ratio}\equiv\frac{\Delta_{r}R}{\Delta_{e}R}=-\frac{e\Delta r}{r\Delta e}=-\frac{\frac{1}{2}(e_{t}+e_{0})\Delta r}{\frac{1}{2}(r_{t}+r_{0})\Delta e}=-\frac{(e_{t}+e_{0})(r_{t}-r_{0})}{(r_{t}+r_{0})(e_{t}-e_{0})}.$ (11)
As we can calculate efficiency $e_{t}$ and redundancy $r_{t}$ at any time $t$, $\Delta_{e}R$ and $\Delta_{r}R$ can be calculated by
$\Delta_{e}R=(R_{t}-R_{0})\frac{\Delta_{e}R}{\Delta_{r}R+\Delta_{e}R}=(R_{t}-R_{0})\frac{1}{1+\mathrm{ratio}}$ (12)
and
$\Delta_{r}R=(R_{t}-R_{0})\frac{\Delta_{r}R}{\Delta_{r}R+\Delta_{e}R}=(R_{t}-R_{0})\frac{\mathrm{ratio}}{1+\mathrm{ratio}}.$ (13)
2.5 Relative contributions of changes in P, PMI, and PMR
Furthermore, we decompose resilience into three internal components and investigate the contributions of changes in these components to changes in resilience (Liang et al., 2020). The first internal component is the concentration degree of trade flows, defined as a P matrix containing element $p_{ij}$ that is expressed as Eq. (2), where a higher $p_{ij}$ indicates a greater trade flow relative intensity.
The second internal component is the inter-dependency among economies, defined as a point-wise mutual information (PMI) matrix containing element $\mathrm{pmi}_{ij}$ that denotes the degree of inter-dependency between economies $i$ and $j$ (Kharrazi and Fath, 2016). A higher value indicates a higher probability of a flow between given economies. The element $\mathrm{pmi}_{ij}$ of PMI matrix is expressed as:
$\mathrm{pmi}_{ij}=\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=\ln\frac{f_{ij}s}{s_{i}^{\mathrm{out}}s_{j}^{\mathrm{in}}}.$ (14)
According to Eq. (1), we start by calculating contributions of changes in $p_{ij}$ and $\mathrm{pmi}_{ij}$ to changes in efficiency by the index decomposition method (Ang and Zhang, 2000):
[…]
where $\Delta e_{ij}$ denotes changes in efficiency caused by changes in the trade link from economy $i$ to $j$. Hence, the contributions of changes in P and PMI to changes in efficiency can be written as
$\Delta e=\Delta_{p}e+\Delta_{\mathrm{pmi}}e,$ (16)
where
$\Delta_{p}e=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta_{p}e_{ij}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta p_{ij}\times\frac{1}{2}\left(\mathrm{pmi}_{ij}^{0}+\mathrm{pmi}_{ij}^{t}\right)$ (17)
and
$\Delta_{\mathrm{pmi}}e=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta_{\mathrm{pmi}}e_{ij}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta\mathrm{pmi}_{ij}\times\frac{1}{2}\left(p_{ij}^{0}+p_{ij}^{t}\right)$ (18)
denote the contributions of changes in the concentration degree of trade flows and those in the inter-dependency among economies to changes in network efficiency, respectively.
The third internal component is the inter-independency among economies, defined as a point-wise mutual redundancy (PMR) matrix containing element $\mathrm{pmr}_{ij}$ that denotes the degree of inter-independency between economies $i$ and $j$ (Liang et al., 2020). A higher value indicates a higher diversity for the origin and destination of flows between given economies. Similarly, the element $\mathrm{pmr}_{ij}$ of PMR matrix is expressed as
$\mathrm{pmr}_{ij}=-\ln p_{j\mid i}-\ln p_{i\mid j}=\ln\frac{s_{i}^{\mathrm{out}}s_{j}^{\mathrm{in}}}{f_{ij}^{2}}.$ (19)
Hence, the contributions of changes in P and PMR to changes in redundancy can be written as:
$\Delta r=\Delta_{p}r+\Delta_{\mathrm{pmr}}r,$ (20)
where
$\Delta_{p}r=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta_{p}r_{ij}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta p_{ij}\times\frac{1}{2}\left(\mathrm{pmr}_{ij}^{0}+\mathrm{pmr}_{ij}^{t}\right)$ (21)
and
$\Delta_{\mathrm{pmr}}r=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta_{\mathrm{pmr}}r_{ij}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\Delta\mathrm{pmr}_{ij}\times\frac{1}{2}\left(p_{ij}^{0}+p_{ij}^{t}\right)$ (22)
denote the contributions of changes in the concentration degree of trade flows and those in the inter-independency among economies to changes in network redundancy, respectively.
Subsequently, we can rewrite Eqs. (12) and (13) as
$\Delta_{e}R=\mathrm{peR}\times\Delta e=\mathrm{peR}\times(\Delta_{p}e+\Delta_{\mathrm{pmi}}e)$ (23)
and
$\Delta_{r}R=\mathrm{prR}\times\Delta r=\mathrm{prR}\times(\Delta_{p}r+\Delta_{\mathrm{pmr}}r),$ (24)
where peR and prR denote changes in resilience caused by unitary changes in efficiency and redundancy, expressed respectively as
$\mathrm{peR}=\frac{\Delta_{e}R}{\Delta e}=\frac{R_{t}-R_{0}}{e_{t}-e_{0}}\times\frac{1}{1+\mathrm{ratio}}$ (25)
and
$\mathrm{prR}=\frac{\Delta_{r}R}{\Delta r}=\frac{R_{t}-R_{0}}{r_{t}-r_{0}}\times\frac{\mathrm{ratio}}{1+\mathrm{ratio}}.$ (26)
Ultimately, we decompose changes in resilience into three components:
[…]
where $\Delta_{p}R$, $\Delta_{\mathrm{pmi}}R$, and $\Delta_{\mathrm{pmr}}R$ denote the contributions of changes in P, PMI, and PMR to changes in resilience, respectively.
2.6 Significance of economies and trade relationships in resilience
The perturbation methods, including node removal, link removal, and the global weight perturbation, are commonly employed to investigate changes in system resilience (Liu et al., 2024a). Since the global weight perturbation does not affect results in Eqs. (1) and (5), according to the leave-one-out approach (Hué et al., 2019; Zedda and Cannas, 2020), we quantify the contributions of changes in economies and relationships to changes in resilience to identify critical economies and trade relationships from the resilience perspective.
We denote the resilience of the networked food trade system $G$ with $N_{\mathcal{V}}$ economies and $N_{\mathcal{E}}$ trade relationships as $R^{G}$. For a given economy $i$, we consider a system $G_{(-i)}$ with $N_{\mathcal{V}}-1$ nodes that includes all economies except economy $i$ and denote the resilience of this system as $R^{G_{(-i)}}$. Hence, the contributions of changes in economy $i$ can be expressed as
$\Delta R_{i}=\frac{\lvert R^{G}-R^{G_{(-i)}}\rvert}{R^{G}},$ (28)
which also denotes the significance of economy $i$ in the international food trade system from a network resilience perspective. A higher $\Delta R_{i}$ implies that economy $i$ plays a more important role in system resilience.
Similarly, for a given trade relationship $e_{ij}$, we consider a system $G_{(-e_{ij})}$ with $N_{\mathcal{E}}-1$ links that includes all relationships except the relationship $e_{ij}$ and denote the resilience of this system as $R^{G_{(-e_{ij})}}$. Hence, the contributions of changes in relationship $e_{ij}$ can be expressed as
$\Delta R_{e_{ij}}=\frac{\lvert R^{G}-R^{G_{(-e_{ij})}}\rvert}{R^{G}},$ (29)
which also denotes the significance of relationship $e_{ij}$ in the international food trade system from a network resilience perspective. A higher $\Delta R_{e_{ij}}$ implies that relationship $e_{ij}$ plays a more important role in system resilience.
2.7 Resilience of core and peripheral food trade networks
The international food trade network exhibits a pronounced core-periphery structure (Dong et al., 2018), which prompts us to investigate the differences between resilience of core and peripheral food trade networks. It should be emphasized that our study explicitly excludes cumulative economy removal scenarios, as termination of trade relationships remains operationally feasible (Guo et al., 2025), which fundamentally differs from complete embargoes on economy’s food trade. We sort all trade relationships either in ascending or descending order of $\Delta R_{e_{ij}}$ and define $p=\frac{m}{N_{\mathcal{E}}}$ as the ratio of the number of relationships within the ordered set $\mathcal{E}^{\prime}$ to the total number of relationships $N_{\mathcal{E}}$, where the set of given trade relationships =$\mathcal{E}^{\prime}=\{e_{1},e_{2},\dots,e_{m}\}$ and $1\leq m\leq N_{\mathcal{E}}$.
Hence, $G_{(-p)}$ is defined as the food trade network that includes the proportion $1-p$ of relationships in order of $\Delta R_{e_{ij}}$. For ascending order of $\Delta R_{e_{ij}}$, the network $G_{(-p)}$ can be regarded as a core food trade network when the value of $p$ is large (that is, the network is constituted by relationships with greater impact on resilience). Also, for descending order of $\Delta R_{e_{ij}}$, the network $G_{(-p)}$ can be regarded as a peripheral food trade network when the value of $p$ is large (that is, the network is constituted by relationships with less impact on resilience). Subsequently, we can calculate the resilience of core or peripheral food trade networks $R^{G_{(-p)}}$ with respect to the changes in $p$, which may help clarify the role of the above networks in sustaining network resilience and guide structure optimization for the international food trade.
3 Dynamic evolution of the international food trade network and its resilience
Figure 1: The evolution of topological properties of international food trade networks. a,b, The numbers of nodes $N_{\mathcal{V}}$ and links $N_{\mathcal{E}}$ are two alternative measures of network size. c,d, Average degree $\langle k\rangle$ and average strength $\langle s\rangle$ represent the average number of trade partners and trade volume of economies, respectively. e, Average clustering coefficient $\langle C\rangle$ represents the average actual connections among the neighbors of individual nodes. f, Density $\rho$ reflects the level of interconnectedness among economies. g, Average shortest path length $L$ represents the level of separation among economies. h, Diameter $D$ represents the maximum of shortest path lengths between any two economies. i, Modularity $Q$ represents the denseness of flows within a community compared to between communities.
Given that maize, rice, soybean, and wheat are foods that account for most of the calories consumed by the world’s population (D’Odorico et al., 2014; Zhang and Zhou, 2023), we select the international bilateral trade quantities of maize (CPC code [^1] Central Product Classification (CPC) code, a product classification for goods and services promulgated by the United Nations Statistics Division.: 0112), rice (CPC code: F0030), soybean (CPC code: 0141), and wheat (CPC code: 0111) from 1986 to 2022, provided by the “detailed trade matrix” dataset [^2] FAOSTAT, www.fao.org/faostat/en/#data/TM. of the FAO to construct networks, where economies are considered as nodes and trade relationships between economies serve as directed and weighted links with the trade quantities being the weights of links. We note that the abbreviations for economies used in the figures and tables are taken from the International Standard Organization [^3] ISO, www.iso.org/iso-3166-country-codes.html..
Generally, there is a trend of increasing network size (Fig. 1a,b). The size of rice trade network is the largest, followed by that of the maize, wheat, and soybean trade networks. However, the average strength for rice is always lower than that of other foods due to its low trade volume (Fig. 1d). In 2022, for example, the total quantities traded globally for maize, soybean, and wheat are 180 million tons, 168 million tons and 217 million tons, respectively, compared to only 55 million tons for rice. Different from other foods with industrial production attributes, rice is predominantly edible and has a homogenous consumption structure. Also, high self-sufficiency on the rice-producing side and entrenched regional consumption habits [^4] For example, China imports only 2% of its rice consumption in broken rice (Valera et al., 2024). result in low trade volume for rice (Lakkakula et al., 2015). Other reasons are analyzed later. It is worth noting that the number of economies participating in global trade has increased significantly in 2014 with some African and Oceanian economies joining trade networks, driven by a surge in demand from other developing economies. For example, Papua New Guinea and Mozambique, relying on resources of undeveloped arable land, have for the first time exported foods such as maize to large importers [^5] OECD-FAO Agricultural Outlook 2013-2022, https://openknowledge.fao.org/handle/20.500.14283/ar251e..
Figure 2: The international trade networks of maize, rice, soybean, and wheat in 2022. a-d, The international trade network of maize $(N_{\mathcal{V}}=185,N_{\mathcal{E}}=2531)$, rice $({N_{\mathcal{V}}}=173,N_{\mathcal{E}}=3130)$, soybean $({N_{\mathcal{V}}}=171,N_{\mathcal{E}}=1453)$, and wheat $({N_{\mathcal{V}}}=175,N_{\mathcal{E}}=1703)$ in 2022. e-h, The logarithmic-transformed boxplots of four indicators in 2022, including in-degree ($k^{\mathrm{in}}$), out-degree ($k^{\mathrm{out}}$), in-strength ($s^{\mathrm{in}}$), and out-strength ($s^{\mathrm{out}}$).
Meanwhile, we find that the clustering coefficients of the international food trade networks are increasing (Fig. 1e) and the average shortest pathway length of the international access to foods fluctuates markedly (Fig. 1g), especially during food crises (e.g., 07/08 and 11/12) and around major events (e.g., the dissolution of the Soviet Union in 1991 and the COVID-19 pandemic since 2020). In addition, since relatively large modularity values between 0.3 and 0.7 are considered an indicator of a well-structured network partition (Newman, 2004), we find that flows within communities are denser compared to flows between communities for maize, rice, and wheat (Fig. 1i).
Specifically, we analyze the international trade networks of four foods in 2022, as shown in Fig. 2. The node size refers to its total trade volume (including imports and exports) and we label the top 10 economies by the size (Fig. 2a-d). Compared to maize and rice, the soybean and wheat trade networks exhibit sparser connectivity and lower density (Fig. 1f), manifested by comparable numbers of participating economies but fewer trade relationships (Fig. 1a,b). For both of them, the first quartile is zero in the logarithmic-transformed boxplots of $k^{\mathrm{out}}$ and $s^{\mathrm{out}}$ (Fig. 2f,h), indicating their more concentrated exporting characteristics. The difference, however, is that trade volume of wheat varies less among importers and is more evenly distributed (Fig. 2g). Moreover, the soybean trade network exhibits substantial heterogeneity in the size of participating economies, indicating highly monopolistic characteristics by a few economies such as the United States, Brazil, and China (Fig. 2c).
Overall, evidently, the United States and China are the most important trading economies, with the former being a major exporter and the latter a major importer. With regard to exports, the United States has vast arable land and suitable climatic condition for the large-scale cultivation of a wide range of foods. Not only in production but also in variety, the adequate level of supply has also contributed to it being a major exporter. With regard to imports, although China has a large area of arable land, it has a large population and a small per capita area of arable land, which causes China having to rely on imports to meet domestic demands.
Figure 3: The evolution of several properties of international food trade networks. The resilience (a), efficiency (b), redundancy (c), and $\alpha$ values (d) of international food trade networks from 1986 to 2022. In this figure, the red dotted line indicates $1/{\mathrm{e}}\approx 0.3679$. e,f, The correlations between resilience and two properties: efficiency and redundancy.
Furthermore, we calculate efficiency, redundancy, $\alpha$ values, and resilience of the international food trade networks from 1986 to 2022 (Fig. 3a-d). Overall, trade networks of maize, rice, and wheat are more resilient. Separately, maize and wheat exhibit parallel enhancements in both efficiency and redundancy, resulting in the better trade-off within $\alpha$ framework. Rice exhibits enhancement in redundancy but fluctuation in efficiency, resulting in a slowly declining resilience in recent years. Conversely, the soybean trade network is not only more vulnerable, its resilience fluctuates more significantly. As is known, approximately 87% of soybean export quantities originated from Brazil and the United States in 2022. Due to the excessively high concentration of exporters, most economies, except for the main exporters, remain relatively isolated, with limited alternative trade pathways. Consequently, the soybean trade network lacks both efficiency and redundancy, and exhibits an insufficient trade-off within the $\alpha$ framework, ultimately resulting in lower resilience. Notably, $\alpha$ values of all networks are less than $1/{\mathrm{e}}$, resulting in the highly positive correlation between efficiency and resilience (Fig. 3e) and the negative correlation between redundancy and resilience (Fig. 3f). It indicates the current trade networks possess sufficient redundancy but require enhanced efficiency, especially for soybean.
4 Network properties and resilience of the international food trade network
4.1 Correlations between topological properties and resilience
We investigate the correlations between topological properties of the international food trade networks and their resilience, as shown in Fig. 4. Regardless of foods, evidently, we find that modularity exhibits the strongest positive correlation to resilience, indicating that more resilient states of international trade may be associated with stronger community structures (Garlaschelli and Loffredo, 2005; Li et al., 2024). A higher modularity reflects denser intra-community flows relative to inter-community flows, corresponding to processes of regional integration that leads to stronger inter-dependence among economies within the community and enhanced overall network efficiency (de Raymond et al., 2021). For maize and wheat, for example, the correlation coefficients between resilience and modularity are 0.72 and 0.80, respectively.
Figure 4: The correlations between topological properties of the international food trade networks and their resilience. a-i, Scattered distribution of resilience and topological indicators. In this figure, the red dotted line indicates $1/{\mathrm{e}}$. j, Kendall’s correlation coefficients between resilience and topological indicators. In this figure, $R_{m}$, $R_{r}$, $R_{s}$, and $R_{w}$ denote resilience of trade networks of maize, rice, soybean, and wheat, respectively.
The correlation between resilience and other topological properties varies across different foods. For example, the network size and clustering characteristics are positively correlated with resilience for maize, soybean, and wheat, but are negatively correlated with resilience for rice. In mature trade networks such as those for maize and wheat, evident clustering characteristics enhance trade efficiency (e.g., through regional alliances and international cooperation organizations, such as NAFTA [^6] NAFTA, The North American Free Trade Area. and MERCOSUR [^7] MERCOSUR, The Southern Common Market.), while the larger network size increases systemic redundancy. Together, these features jointly contribute to sustaining network resilience. This finding fully supports the view that internationalization does not undermine, but rather strengthens the resilience of mature food trade networks (Sartori and Schiavo, 2015).
However, resilience of the rice trade network appears to decline with increases in both clustering coefficients and the network size, potentially due to its inherent network characteristics and the nature of rice as a commodity. The network is heavily concentrated around a few major exporters (such as India, Thailand, and Vietnam) and importers (such as China and the Philippines), forming a single-core structure of Asia. Despite greater connectivity, this centralized architecture has exacerbated systemic vulnerability during network expansion. Meanwhile, while geographic proximity suggests potential trade alliances, the lack of strategic complementarity, exemplified by competitive underpricing between Vietnam and Thailand, undermines cluster resilience by fostering internal fragmentation. Additionally, as an inelastic staple commodity with low substitution elasticity, the rice trade network faces catastrophic systemic collapse once major Asian producers implement export restrictions to ensure domestic food security (e.g., India’s 2022 export ban for rice).
4.2 Relative contributions of changes in efficiency and redundancy to changes in resilience
We calculate the relative contributions of changes in efficiency and redundancy to changes in network resilience for 1986-2022 (Fig. 5a-d). As previously mentioned, the current trade networks are relatively redundant due to $\alpha<1/{\mathrm{e}}$. Hence, changes in resilience has been dominated by changes in efficiency. For example, although the redundancy effects in the soybean trade network are lower than those in the base year prior to 2004, the network resilience is also higher during this period because of continued positive changes in efficiency (Fig. 5c). The exception is rice. Before 1998, changes in resilience of the rice trade network has been dominated by changes in efficiency. However, the effect has been diverse after 2010 with the dominance of redundancy (Fig. 5b), and the reason (increasing redundancy and oscillating efficiency) can also be found in Fig. 3b.
Figure 5: Relative contributions of changes in efficiency and redundancy to changes in resilience of international food trade networks for 1986-2022 (a-d) and specific time periods (e-h).
We also calculate the relative contributions of changes in efficiency and redundancy to changes in network resilience for specific time periods (Fig. 5e-h). Different from what is found above, we find that changes in resilience are not always dominated by one property or another. For example, the 0.03 increase (from 0.19 to 0.22) in resilience of the soybean trade network during 1986 to 1992 is due to the 128% increase in efficiency, but the 0.006 increase (from 0.272 to 0.278) in resilience of that during 2004 to 2010 is due to the 363% increase in redundancy (Fig. 5g). Moreover, the 0.02 increase (from 0.32 to 0.34) in resilience of the wheat trade network during 1992 to 1998 is due to the 127% increase in efficiency, but the 0.02 decrease (from 0.342 to 0.34) in resilience of the wheat trade network during 1998 to 2004 is due to the 318% decrease in redundancy (Fig. 5h). The findings suggest that the dominance of changes in efficiency to changes in resilience is diminishing while $\alpha$ values are closer to $1/{\mathrm{e}}$.
4.3 Relative contributions of changes in P, PMI, and PMR to changes in resilience
We calculate the relative contributions of changes in the concentration degree of trade flows, economies inter-dependency, and economies inter-independency to changes in network resilience for 1986-2022 (Fig. 6a-d). Overall, changes in the concentration degree of trade flows contribute more to changes in resilience for maize, soybean, and wheat throughout the entire period of 1986-2022, indicating the greater influence of trade flows concentration for food demand on resilience (Liang et al., 2020). However, these effects are diminishing with trade globalization, resulting in negative impact on changes in resilience. Indeed, these findings reflect the underlying shift in food trade discourse from flow monopolies to strategic interactions within relational networks. For example, resilience for rice has been dominated by changes in the concentration degree of trade flows until 2010. Since then, effects of relationships among economies has been highlighted, resulting in decreasing resilience. Although the effect for the concentration degree of trade flows of the rice trade network is higher than 5% after 2010, changes in resilience exhibit phase inversion due to synergistic interaction for interrelationships (Fig. 6b).
Figure 6: Relative contributions of changes in the concentration degree of trade flows (P), economies inter-dependency (PMI), and economies inter-independency (PMR) to changes in resilience of international food trade networks for 1986-2022 (a-d) and specific time periods (e-h).
We also calculate the relative contributions of changes in the concentration degree of trade flows, economies inter-dependency, and economies inter-independency to changes in network resilience for specific time periods (Fig. 6e-h). We find that changes in resilience are dominated by the concentration degree of trade flows and economies inter-dependency, corroborating prior observations that efficiency-driven mechanisms exert predominant influence. Exceptionally, the 0.001 decrease (from 0.262 to 0.261) in resilience of the soybean trade network during 2016 to 2022 is due to the 620% decrease in economies inter-independency (Fig. 6g).
5 Identifying critical economies and trade relationships from the resilience perspective
5.1 Relative contributions of changes in economies and trade relationships to changes in resilience
To quantify the relative contributions of changes in economies and trade relationships to changes in resilience, following the leave-one-out approach, we remove some given economy or trade relationship independently and regard changes in the network resilience as the significance of the economy or trade relationship. Figs. 7 and 8 depict the correlation between trade volume of the economy and trade relationship and their contributions to resilience during 1992 to 2022, respectively.
Figure 7: The correlations between trade volume of economy $i$ and changes in resilience of the international food trade network as the economy $i$ is removed during 1992 to 2022. a-f, The maize trade networks. g-l, The rice trade networks. m-r, The soybean trade networks. s-x, The wheat trade networks.
Figure 8: The correlations between trade volume from economy $i$ to $j$ and changes in resilience of the international food trade network as the relationship $e_{ij}$ is removed during 1992 to 2022. a-f, The maize trade networks. g-l, The rice trade networks. m-r, The soybean trade networks. s-x, The wheat trade networks.
Overall, the scale-free characteristics of the international food trade system are reaffirmed from the perspective of changes in resilience caused by economies and trade relationships. That is, a small number of economies and trade relationships contribute most of the changes in resilience, and their influences on resilience demonstrates a power-law distribution. For example, in 1998, only four economies and a single trade relationship contribute more than 2% to resilience in the international wheat trade network, while all remaining economies and relationships each contributes less than 1% (Fig. 7t and Fig. 8t). Over time, the disparity in participants’ contributions to network resilience has narrowed, particularly in the maize and wheat trade networks, indicating that mature networks possess more balanced structures. For example, while four economies and a single trade relationship contribute more than 2.5% to resilience of the international maize trade network in 1992 and 1998 (Fig. 7a,b and Fig. 8a,b), only a single economy contributes more than 2% to that in 2022 (Fig. 7f), as well as all trade relationships contributing less than 1% (Fig. 8f). This finding suggests that more economies and trade relationships are integrating in the system and lessening the significance of the dominators and thus making the system more balanced and resilient, aligning with previous findings (Sartori and Schiavo, 2015).
To further investigate the heterogeneity among leading contributors, we respectively present the top 10 economies and trade relationships in 2022 in Tables LABEL:Table_Economies and LABEL:Table_Relationships, based on their relative contributions to resilience.
Overall, the contributions of leading economies and trade relationships to resilience are highest in the soybean trade network, followed by those in rice, maize, and wheat, which aligns with the resilience characteristics discussed above. It shows that the top 10 economies’ $\Delta R$ values range from 1.64% to 17.06% for soybean, compared to 0.82%-5.79% for rice, 0.62%-2.30% for maize, and 0.34%-1.14% for wheat. Also, the top 10 trade relationships’ $\Delta R$ values range from 1.75% to 23.59% for soybean, 0.70%-2.55% for rice, 0.49%-0.74% for maize, and 0.23%-0.31% for wheat. These findings suggest that the soybean trade network exhibits substantially greater vulnerability to disruptions compared to its maize and wheat counterparts.
In terms of economies, we find that with the exception of China, which is the main importer, the top 3 contributors are all exporters, including Brazil, Ukraine, the United States, India, Thailand, Canada, Russia, and France. This underscores the greater influence of an economy’s export role on network resilience (Grassia et al., 2022). Generally, export restrictions raise the risk of supply chain disruptions and trigger cascading effects that undermine the resilience (Burkholz and Schweitzer, 2019). For example, maize exports from Ukraine have been disrupted by the Russia-Ukraine conflict, which not only destabilizes maize imports in other economies but also further shocks the supply of pork and poultry meat due to the lack of fodder maize (Laber et al., 2023). Meanwhile, export restrictions tend to disrupt the global food price and the resulting price volatility induces the trade resilience. For example, after Ukraine’s food exports are disrupted, 18 economies impose food export bans, leading to a 23% increase in the global food price index. In contrast, import restrictions primarily impact unilateral market access, whereas exporters can mitigate risks through trade partner diversification and inventory expansion, thereby exerting only a limited impact on overall trade resilience.
Comparison reveals that the contributions of leading economies to network resilience are generally higher than those of primary trade relationships, which can be attributed to the latter’s higher redundancy. For example, while Japan’s maize imports from the United States constitute a critical trade linkage, supply chain disruption in this corridor induces merely a 0.74% impact on the network resilience, owing to substitutable sourcing from Brazil. Reciprocally, Brazilian maize export restrictions would induce a 2.30% impact on the network resilience. The soybean trade network constitutes a notable exception, where individual economy resilience contributions from Brazil (17.06%), China (16.89%), and the United States (13.44%) are eclipsed by the bilateral trade linkages: the United States-China (23.59%) and Brazil-China (20.95%), demonstrating critical path inter-dependency and lack of redundancy.
Although trade volume is significantly positively correlated with the resilience contributions of economies and trade relationships, the two should not be directly equated, particularly when evaluated through the lens of resilience. For a comprehensive comparison, we also present the rankings of economies based on degree centrality and betweenness centrality, two of the most commonly used metrics in the network science literature (Li et al., 2013; Lü et al., 2016; Lambiotte et al., 2019; Liu et al., 2024b), in Table LABEL:Table_Economies. Degree centrality captures the importance of an economy in terms of the number of its trade partners, whereas betweenness centrality reflects its role as a bridge in the network, measured by the number of shortest paths that pass through it. We find that trade volume, degree centrality, and betweenness centrality rankings more accurately capture the strategic importance of economies such as the United States, Germany, France, and Canada. However, all three metrics are fundamentally volume-based and may overlook the dynamic topological structures (Qi and Mei, 2024) and the inter-dependence between nodes (He et al., 2024) within the network. Consequently, these metrics may fail to capture the significance of certain economies such as Paraguay and Kazakhstan with moderate trade volume or low network centrality.
For example, Paraguay ranks sixth (0.75%), eighth (1.46%), and fifth (7.13%) in resilience contributions to the maize, rice, and soybean trade networks, respectively, despite falling outside the top 20 in trade volume and the top 50 in both degree and betweenness centrality for all three commodities. Three potential explanations may account for this phenomenon. First, serving as a strategic redundancy reserve in South America’s multimodal corridors, Paraguay leverages its geopolitical advantage as a landlocked state to activate the Paraguay-Paraná Waterway intermodal system during congestion mitigation protocols at primary Brazilian ports (e.g., Santos). This enables maize rerouting through alternative terminals like Paranaguá, sustaining 15% of displaced shipments while maintaining supply chain continuity. Second, to address the demands of Brazil’s premium culinary market, Paraguay operationalizes a “low-volume and high-premium” trade paradigm by importing Thai Jasmine rice for value-added processing before re-exporting to Brazil. This strategic intervention mitigates production shortfalls in domestic indica rice cultivation. It is further substantiated by the elevated $\Delta R$ values in Paraguay-Brazil and Paraguay-Argentina, as quantified in Table LABEL:Table_Relationships. Third, Paraguay stands as the sole nation maintaining 100% Non-GMO [^8] Non-GMO, Products that are produced without genetic engineering. soybean commercial cultivation globally, positioning itself to fulfill the European Union’s stringent Non-GMO demand (evidenced by a 58% export surge in 2022), thereby securing an irreplaceable supply chain role. Similarly, Kazakhstan ranks fourth (0.64%) in resilience contribution to the wheat trade network, despite falling outside the top 20 in trade volume, degree centrality, and betweenness centrality for this commodity. Its critical role stems from functioning as a crisis replacement corridor, offering alternative routes, such as the China-Europe Railway and the Trans-Caspian Corridor, when the Black Sea passage or conventional land routes are disrupted. Moreover, Kazakhstan supplies 65% of China’s organic wheat imports for animal feed, fulfilling the demand for “Non-GMO and high-gluten” products, thereby securing an indispensable position in the international wheat trade system. Hence, compared to these commonly used methods, we argue that our approach effectively identifies important economies and relationships that possess low volumes but indispensable roles from a resilience perspective. Meanwhile, cases such as Paraguay and Kazakhstan demonstrate that strategic trade positioning may surpass the indiscriminate pursuit of trade volume in enhancing contributions to the international food trade resilience, which may be an indication for other economies.
5.2 Comparing resilience of core and peripheral food trade networks
After calculating the relative contributions of trade relationships to resilience, we sort the results in ascending or descending order and investigate the evolution of resilience of core or peripheral networks as trade relationships are removed cumulatively, as shown in Figs. 9 and 10, respectively.
Figure 9: The evolution of efficiency, redundancy, and resilience of the international food trade sub-networks as trade relationships are removed cumulatively in descending order of $\Delta R$. The red dotted line indicates $1/{\mathrm{e}}$. a-c, The maize trade networks. d-f, The rice trade networks. g-i, The soybean trade networks. j-l, The wheat trade networks.
Table 1: The efficiency, redundancy, and resilience of the optimal international food trade sub-network after $p$ proportion relationships are removed cumulatively in descending order of $\Delta R$.
1992 1998 2004 2010 2016 2022
Panel A: Maize
$p$ 5.95% 5.72% 3.76% 4.48% 7.83% 4.78%
Efficiency 1.7383 1.8795 1.9145 1.9925 2.1229 2.0733
Redundancy 2.9337 3.2574 3.2902 3.4206 3.6481 3.5704
Resilience 0.3679 0.3679 0.3679 0.3679 0.3679 0.3679
Panel B: Rice
$p$ 27.42% 26.32% 35.13% 40.16% 45.11% 43.37%
Efficiency 2.0214 2.1725 2.2976 2.3452 2.4174 2.3843
Redundancy 3.4729 3.7278 3.9480 4.0295 4.1532 4.0953
Resilience 0.3679 0.3679 0.3679 0.3679 0.3679 0.3679
Panel C: Soybean
$p$ 15.04% 14.15% 11.71% 8.72% 6.47% 5.86%
Efficiency 1.6021 1.6893 1.7729 1.8033 1.8719 1.9211
Redundancy 2.7404 2.9093 3.0454 3.1039 3.2028 3.2823
Resilience 0.3679 0.3679 0.3679 0.3679 0.3679 0.3679
Panel D: Wheat
$p$ 26.91% 29.17% 26.45% 24.25% 22.97% 21.15%
Efficiency 1.8343 1.9793 2.1144 2.1347 2.1696 2.1475
Redundancy 3.1570 3.4002 3.6362 3.6681 3.7284 3.6954
Resilience 0.3679 0.3679 0.3679 0.3679 0.3679 0.3679
Notes: $p$ refers to the proportion of the number of the removed links to the total number of links.
Figure 10: The evolution of efficiency, redundancy, and resilience of the international food trade sub-networks as trade relationships are removed cumulatively in ascending order of $\Delta R$. The red dotted line indicates $1/{\mathrm{e}}$. a-c, The maize trade networks. d-f, The rice trade networks. g-i, The soybean trade networks. j-l, The wheat trade networks.
Fig. 9 reveals that the observed resilience enhancement directly correlates with efficiency gains, corresponding to the lack of efficiency caused by core trade relationships, aligning with previous findings (Karakoc and Konar, 2021). In other words, overconcentration of flows along core trade pathways can, in fact, reduce both efficiency and resilience of the food trade network. As $p$ increases, resilience of the trade network for maize, rice, soybean, or wheat rises to reach the maximum value ($1/{\mathrm{e}}$) before declining, suggesting that there exists at least one sub-network with optimized resilience, and its metrics are shown in Table 1. Although maize and soybean networks exhibit similar optimal $p$ values, their efficiency effects differ fundamentally. In 2022, for example, the trade-off between efficiency and redundancy in the maize trade network is due to the 68% increase (from 1.23 to 2.07) in efficiency after core relationships are removed. However, the trade-off in the soybean trade network requires the 419% surge (from 0.37 to 1.92) in efficiency, revealing acute oligopolistic vulnerability where the top 2 relationships contribute more than 40% to the network resilience. Also, food trade networks exhibit sharp decline in resilience as extensive trade relationships are removed, while the phenomenon’s $p$ value tends to be closer to 1 in recent years. For example, resilience of the maize trade network decreases sharply (defined as $\Delta R_{p}\leq-\Delta p$ at $p$) after $p=61.41\%$ trade relationships are removed in 1992, while $p=70.43\%$ and $p=91.30\%$ in 2010 and 2022, respectively (Fig. 9c). This finding suggests that the capacity to sustain resilience of peripheral relationships is becoming greater and stronger.
In Fig. 10, conversely, efficiency, redundancy, and resilience remain stable until removing about 75% peripheral trade relationships, consistently confirming that core trading relationships overwhelmingly dominate the international food trade (Wood et al., 2018). However, not all networks formed solely by core trade relationships attain optimal resilience, especially for soybean (Table 2). For instance, the 1992 soybean trade system composed of 84.96% non-core relationships is more resilient than that composed of 14.04% core relationships (0.3679 and 0.2260). This finding highlights the necessary roles of peripheral economies and trade relationships in sustaining international trade network resilience again, through complementary mechanisms such as improving trade efficiency and facilitating risk diversification via diversified trade pathways, indicating the irrationality of hegemonism and the importance of global cooperation.
Table 2: The efficiency, redundancy, and resilience of the optimal international food trade sub-network after $p$ proportion relationships are removed cumulatively in ascending order of $\Delta R$.
1992 1998 2004 2010 2016 2022
Panel A: Maize
$p$ 99.52% 99.20% 98.60% 98.83% 99.53% 98.97%
Efficiency 0.4819 0.5821 0.7778 0.8902 0.7640 1.1172
Redundancy 0.7326 1.2877 1.7081 1.5232 1.5169 1.9101
Resilience 0.3668 0.3633 0.3635 0.3679 0.3664 0.3679
Panel B: Rice
$p$ 94.08% 96.79% 97.95% 97.22% 99.81% 99.74%
Efficiency 1.3683 1.2987 1.2759 1.3855 0.6931 0.7627
Redundancy 2.3503 2.2403 2.2038 2.3903 1.2021 1.3193
Resilience 0.3679 0.3679 0.3679 0.3679 0.3679 0.3679
Panel C: Soybean
$p$ 85.96% 97.43% 99.08% 95.20% 69.33% 98.48%
Efficiency 0.3219 0.3551 0.4829 0.3804 0.3850 0.2883
Redundancy 3.0044 2.2313 1.4116 2.1590 2.6500 1.7528
Resilience 0.2260 0.2726 0.3484 0.2844 0.2619 0.2764
Panel D: Wheat
$p$ 97.67% 94.57% 96.95% 87.57% 95.26% 92.48%
Efficiency 0.8443 1.2918 1.2356 1.6963 1.5625 1.6603
Redundancy 1.5266 2.2482 2.1281 2.9159 2.6974 2.8544
Resilience 0.3677 0.3679 0.3679 0.3679 0.3679 0.3679
Notes: $p$ refers to the proportion of the number of the removed links to the total number of links.
6 Conclusions
Based on the FAO dataset from 1986 to 2022, we construct international trade networks for maize, rice, soybean, and wheat, where economies are represented as nodes and trade relationships among them as directed and weighted links. To address the need of resilience measurement in the networked system, we employ an entropy-based approach grounded in information theory. Subsequently, we explore the correlations between resilience and topological properties. We also employ the index decomposition analysis to quantify the relative contributions of several properties to resilience, including efficiency, redundancy, trade flows concentration, and interrelationship among economies. Furthermore, we apply the leave-one-out approach to identify critical economies and trade relationships from a resilience perspective. Given the pronounced core-periphery structure of networks, we also discuss the roles of core and peripheral components in sustaining network resilience. Our conclusions are presented below.
6.1 Theoretical contributions
We have several theoretical contributions to the literature on food trade network resilience. First, we introduce a method based on efficiency and redundancy—two fundamental system properties—to measure resilience, drawing from information theory. Our review of existing food trade network studies highlights the need for robust resilience metrics. Unlike prior research that typically relies on changes in unweighted topological properties, we adopt an entropy-based approach originally developed for ecological networks to assess the resilience of the international food trade system. This approach incorporates both the micro perspective of economy and inter-economies behavior and the macro perspective of dynamic network structure, which better focuses on the influence of trade links and flows and effectively measures resilience in food trade systems (addressing RQ1: How can the resilience of the international food trade system can be effectively measured?).
Second, we employ an index decomposition analysis to investigate the contributions of internal components to resilience. Most driving mechanisms in relevant literature focus on topological properties and external shocks, which may overlook the evolution of resilience caused by internal components or structures. To clarify these mechanisms, we decompose resilience into five internal components and investigate their contributions to changes in resilience, including efficiency, redundancy, the concentration degree of trade flows, the inter-dependency among economies, and the inter-independency among economies (addressing RQ2: How do internal components contribute to changes in resilience of the international food trade system?).
Third, by applying a leave-one-out approach, we identify critical economies and trade relationships in the international food trade network from a resilience perspective. While recent literature often employs centrality metrics (e.g., degree centrality and betweenness centrality) to assess the importance of individual economies, probably overlooking critical players with low trade volume or centrality. Moreover, compared to the research on assessing the importance of nodes, that of links are relatively insufficient. Our approach complements the existing literature by capturing critical trade relationships and previously overlooked economies, offering a novel insight compared to traditional metrics (addressing RQ3: How can critical economies and trade relationships in the international food trade system be identified from a resilience perspective?).
6.2 Findings and managerial implications
The results reveal several interesting findings and inspired by these, we summarize several managerial implications.
First, efficiency remains the primary driver of resilience, while current international trade networks tend to exhibit greater redundancy within the $\alpha$ trade-off framework. Furthermore, we find that clearer community partitions can enhance resilience by promoting denser intra-community trade flows and improving overall trade efficiency, as observed in networks for maize, rice, and wheat. At present, therefore, strengthening intra-community trade linkages through regional alliances (e.g., NAFTA, MERCOSUR, G20, ASEAN, and AfCFTA) would greatly contribute to maintaining resilience of trade in these foods (de Raymond et al., 2021), particularly during major disruptions such as the COVID-19 pandemic (Wei et al., 2025). In contrast, the soybean trade network appears more vulnerable due to the highly monopolistic positions of major economies such as the United States, Brazil, and China, resulting in lower modularity and efficiency. Hence, in terms of trade security and economic benefit, economies should strengthen regional production cooperation and technical exchanges to improve soybean production capacity and trade efficiency (Ma et al., 2024). This would not only improve the diversification of trade linkages and risk resistance but also jointly ensure domestic soybean supply and promote industrial development.
Second, rice exhibits distinct trends and characteristics compared to other staple foods, necessitating more targeted attention. We find that several topological properties—such as network size and clustering coefficient—are positively correlated with resilience for maize, soybean, and wheat but not for rice. The resilience of the rice trade network tends to decline with increasing globalization, which may be attributed to its inherent edibility. However, more critical attention should be directed toward the Asia-centered monocentric structure in the rice trade and the frequent export restrictions imposed by major exporters, both of which fundamentally undermine resilience. Hence, expanding production and liberalizing rice exports are essential for improving rice trade resilience. On one hand, expanding rice production outside of Asia is important for enhancing diversification (Puma et al., 2015). This would not only mitigate the risks of supply disruptions caused by events—such as climate extremes, exports bans, and competitive underpricing—in Asia, but also lower rice access costs in other regions and generate income for many developing countries. On the other hand, liberalizing rice exports of major producers, while ensuring sufficient domestic supply, would not only improve overall trade resilience, but also benefit domestic farmers and economy (Valera et al., 2024).
Third, the underlying discourse on food trade has shifted from flow monopolies to strategic interactions within relational networks, placing greater emphasis on the importance of connectivity (Sartori and Schiavo, 2015). Using index decomposition analysis, we find that the concentration effects of trade flows continue to be dominant drivers of changes in resilience with the growing influence of interrelationships among economies in recent years. These effects are further substantiated by the identification of critical contributors to resilience. Economies with moderate trade volumes or low network centrality—such as Paraguay and Kazakhstan—can significantly contribute to system resilience due to their indispensable roles in specific products or strategic geographies. Therefore, for many developing economies, there is an indication that strategic trade positioning may surpass the indiscriminate pursuit of trade volume in enhancing contributions to the international food trade resilience. Moreover, developing and maintaining robust and diversified international food transport systems is essential for improving access to food across various terrains and urban landscapes. Similarly, strengthening food trade infrastructures—such as ports and railways—to ensure international trade connectivity is equally necessary (FAO, 2021).
Fourth, core food trade networks may undermine overall system efficiency, whereas peripheral networks play a critical role in sustaining resilience by enabling risk diversification through alternative trade pathways. Within the $\alpha$ trade-off framework, we find that one of the main factors limiting efficiency in current trade systems is the overconcentration of flows along core trade links, particularly for soybeans. Therefore, guaranteeing food security requires not only strong trade linkages among developed economies but also greater efforts to establish freer trade bridges between developed and undeveloped economies. Such efforts would not only contribute to reducing hunger in developing economies but also help diversify trade risks in developed economies and enhance systemic resilience. As noted by Mujahid and Kalkuhl (2016), all economies are expected to benefit from a free and equitable food trade system that operates without discrimination.
6.3 Limitations and future research directions
Despite the contributions of our study, several limitations warrant further investigation. First, we do not account for food supply and demand dynamics when adopting the leave-one-out approach. In reality, when an economy or trade relationship is disrupted, affected economies typically adjust their trade strategies to rebalance supply and demand, triggering cascading effects that reshape flow distributions across the entire network. It is therefore important to clarify that our approach is designed solely to assess the contributions of economies or trade links to network resilience. Recent studies have employed agent-based models and stress-testing approaches to capture the dynamics of risk propagation in food systems under shock scenarios (Kuhla et al., 2025; Diem et al., 2025), and have further extended the analytical scope to multilayer cascading risks (Naqvi et al., 2020; Laber et al., 2023). Future research could build on these approaches by simulating trade flow redistributions under various shocks, which would provide deeper insights into the resilience of the international food trade system. Second, we do not apply causal inferences in this paper when investigating the interrelations between resilience and other properties. To address this limitation, future studies could apply more advanced methods—such as machine learning—to better identify and quantify the drivers of resilience within the international food trade network. Third, although the entropy-based framework provides a theoretical metric for resilience, its application should be further validated and adapted in socio-economic networks that differ from ecosystems shaped through natural selection (Kharrazi et al., 2013). Future research could investigate whether further autocatalytic effects can emerge within the socio-economic network structure, enabling the replacement of redundant pathways with more efficient alternatives.
Acknowledgment
This work was partly supported by the National Natural Science Foundation of China (72171083) and the Central Universities’ Program for Building World-Class Universities (Disciplines) and Special Development Guidance—Cultural Heritage and Innovation
Disclosure statement
No potential conflict of interest was reported by the authors.
Data availability
Publicly available dataset is analyzed in this study. This data can be found here: https://www.fao.org/faostat.